### v1

@@ 10 Conclusions
The automotive battery pack status, from an automaker’s point of view, is reviewed, and near-future expectations and technological advancements are presented. Different parts of the system, including major regulatory standards, are presented, and near-future development activities are mentioned. Pack technology is advancing at a rapid pace, and it is expected to see more effective, safer, and cheaper units shortly.
@@ 11 List of abbreviations
  2. 2.
    EMC    Electromagnetic Compatibility .
  3. 3.
    BMS    Battery Management System.
  4. 4.
    ASIL   Automotive Safety Integrity Level .
  5. 5.
    LPM    Liters per Minute.
  6. 6.
    MSD    Manual Service Disconnect .
  7. 7.
    UN     United Nation.
@@
  9. 9.
    ECE    Economic Commission for Europe .
  10. 10.
    HARA   Hazard Analysis and Risk Assessmen.
  11. 11.
    SG       Safety Goal.
@@
    FEMA   Failure Mode and Effect Analysis.
@@
    NMC    Lithium Nickel Manganese Cobalt Oxid.
@@
    LCO    Lithium Cobalt Oxide.
@@
    HVIL   High-Voltage Interlock Loop.
@@
    ECU    Electronic Control Unit.

### v2

@@ 10 Conclusions
The automotive battery pack status, from an automaker’s point of view, is reviewed, and near-future expectations and technological advancements are presented. Different parts of the system, including major regulatory standards, are presented, and near-future development activities are mentioned. Pack technology is advancing at a rapid pace, and it is expected to see more effective, safer, and cheaper units shortly. The following points highlight some of the most important ongoing changes and emerging technologies that will drive the next generation of automotive battery systems:
  • Battery technology is evolving rapidly, with advancements in energy density, safety, and cost-effectiveness expected in the near future.
  • Higher voltage systems (700–800V) are gaining adoption, improving efficiency, reducing losses, and enabling faster charging, though they require updates in infrastructure and power electronics.
  • Sodium-based batteries are emerging as an alternative to lithium-ion batteries, offering advantages such as abundant raw materials, lower costs, and reduced environmental impact, though they remain in early development.
  • Rare-earth-free battery chemistries are being explored to reduce reliance on materials like cobalt and lithium, addressing supply chain concerns and sustainability challenges.
  • TNO battery technology is gaining attention, offering high charging and discharging power capabilities, which could significantly reduce EV charging times and improve overall vehicle performance.
  • Battery swapping systems, such as those introduced by NIO, provide a promising alternative to fast charging, enabling quick battery replacements and addressing infrastructure limitations.
  • Battery pack integration is becoming more sophisticated, with wireless communication and modular designs enhancing flexibility, maintenance, and manufacturability.
  • The role of BMS is expanding, incorporating AI, advanced diagnostics, and cybersecurity measures to enhance performance and longevity.
  • Safety remains a critical priority, with stricter regulatory standards governing battery pack design, testing, and operational safety, particularly for thermal runaway prevention and high-voltage handling.
  • Thermal management is essential for high-performance battery systems, with ongoing advancements in cooling solutions for both fast charging and high-power applications.
  • Charging infrastructure must evolve alongside battery technology, as many existing chargers cannot accommodate 800V systems, requiring innovative solutions for compatibility with lower-voltage charging networks.
  • Second-life applications for EV batteries are expanding, including energy storage for the grid, supporting a more sustainable lifecycle and reducing environmental impact.
  • Battery electrification is extending beyond passenger cars, with significant research efforts in heavy-duty transport, buses, trucks, and aviation, each requiring unique design considerations.
  • Aviation electrification presents new challenges, requiring alternative battery architectures, extreme environmental durability, and cost-effective solutions distinct from ground vehicles.
  • The push for sustainable, cost-effective, and high-performance battery solutions will continue driving advancements in materials, design, and infrastructure, shaping the future of electric mobility.
@@ 11 List of abbreviations
    ASIL  Automotive Safety Integrity Level.
  2. 2.
    BMS  Battery Management System.
  3. 3.
    BTMS  Battery Thermal Management System.
  4. 4.
    ECE     Economic Commission for Europe.
  5. 5.
    ECU     Electronic Control Unit.
  6. 6.
    EMC     Electromagnetic Compatibility.
  7. 7.
@@
    FEMA   Failure Mode and Effect Analysis.
  9. 9.
    HARA   Hazard Analysis and Risk Assessment.
  10. 10.
    HVIL   High-Voltage Interlock Loop.
  11. 11.
@@
    LCO    Lithium Cobalt Oxide.
@@
    LPM    Liters per Minute.
@@
    MSD    Manual Service Disconnect.
@@
    NMC    Lithium Nickel Manganese Cobalt Oxide.
@@
    SG       Safety Goal.
  18. 18.
    UN      United Nation.